OpenStudy (anonymous):

In a Chinese game show, 50% of the players are eliminated each day. If the show starts with 1,000 players, to the nearest integer, how many players will be present at the end of the 8th day of play?

OpenStudy (anonymous):

\[f(x)=\frac{1000}{2^x},\]where \(x\) is the number of days.

OpenStudy (anonymous):

Or, heck, just halve that number 8 times.

OpenStudy (anonymous):

7 times, rather.

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

How many at the end of the 1st day?

OpenStudy (anonymous):

500

OpenStudy (anonymous):

Cool. How many at the end of the second day?

OpenStudy (anonymous):

250

OpenStudy (anonymous):

Cool. How many at the end of the third day?

OpenStudy (anonymous):

125

OpenStudy (anonymous):

Keep going until you get the 8th day.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

125/2 = 62.5 62.5/2 = 31.25 etc.